OpenStudy (anonymous):

Consider the following recurrence relation. P(n) = 0 if n = 0 [P(n − 1)]^2 − n if n > 0 Use this recurrence relation to compute P(1), P(2), P(3), and P(4).

OpenStudy (callisto):

\[P(n) =\begin{cases}0 & n = 0\\ [P(n-1)]^2-n & n > 0\end{cases}\] Let's do P(1) Since n = 1 in this case, we will use P(n) = [P(n-1)]^2-n, i.e. P(1) = (P(0)) ^2 - 1 Consider the P(0) on the right, we know that n=0, so, P(0) = 0, then we get P(1) = (P(0)) ^2 - 1 = 0^2 - 1 = ... Can you complete P(1) and try the rest now?

OpenStudy (anonymous):

plug n=1,2,3,4 in p(n)=[P(n-1)]^2-n but first find P(1) and so on

OpenStudy (anonymous):

What I don't understand is how did we start at p(0).

OpenStudy (anonymous):

it is already given P(0)=0

OpenStudy (anonymous):

So P(1) would equal a = -1?

OpenStudy (anonymous):

P(n)=0,for n=0 or P(0)=0

OpenStudy (callisto):

P(1) is equal to -1, yes.

OpenStudy (callisto):

The principle of recurrence is to find the value of a function with a large input by finding the value of the function with a smaller input. For example, in your case,